OpenStudy (anonymous):

Will fan and medal. A piece of iron wire has a diameter of 0.485mm. If iron has a density of 7.86g/cc, how long(in meters) should you cute a piece of wire to obtain 0.0214 moles of iron?

OpenStudy (ramirocruzo):

Sorry for short answer @Shaekitchen

OpenStudy (anonymous):

where does the 56 come from in step 1 ? @RamiroCruzo

OpenStudy (ramirocruzo):

Its the Gram Atomic Mass of Fe i.e. mass of 1 mole

OpenStudy (anonymous):

okay so then why did you put the (3.14)r^2h ? Where does that come from ? @RamiroCruzo

OpenStudy (ramirocruzo):

Its the formula of volume of cylinder.....All wires are cylinder.....Oppsy.....Just did wrong calculations.....Misplaced Dia instead of Radius

OpenStudy (anonymous):

so does that mean the answer is wrong ? @RamiroCruzo

OpenStudy (ramirocruzo):

\[0.0214mol=\frac{ Req. Mass }{ 56 }\] \[Req. Mass=1.1984gm\] \[Vol.×Density=Req.mass\] \[\pi r ^{2} h \times 7.86gm =1.1984\] \[h=82.52cm\]

OpenStudy (ramirocruzo):

:D Sorry..........